heidi bought a minibright headlamp. her headlamp requires 3 batteries for every 6 hours of use.she is packing 12 batteries for her trip to marengo cave.

one the line below, write a proportion that can be used to find the number of hours (h) heidi can spend exploring the cave.


proportion_____________________


now use your proportion to find the number of hours heidi can spend in the cave.write your answer on the line below.
SHOW ALL WORK





answer___________________hours

Answer by checkley77(12844)

3/12=6/x
cross multiply:
3x=6*12
3x=72
x=72/3
x=24 hours worth of batteries.